Given f(x)= -5x -1, find f(a-1).
Answer by praseenakos@yahoo.com(507)

QUESTION:

Given f(x)= -5x -1, find f(a-1).


ANSWER:

Given f(x)= -5x -1,

To find f(a-1)....replace x by ( a-1) in the given function.

==> f(a-1) = -5(a-1) -1


==> = -5*a -5*-1 -1


==> = -5a + 5 - 1


==> = -5a + 4


So f(a-1) = -5a + 4



Hope you understood.


Regards.

Praseena.